# Elliptic functions, Floquet transform and Bergman spaces on doubly periodic domains

### v1

@@ Abstract.
Given a domain $\Omega$ in the complex plane $\mathbb{C}$}, we denote by $L^{2}(\Omega)$ the usual Lebesgue-Hilbert space with respect to the (real) area measure dAdA and by $A^{2}(\Omega)$ the corresponding Bergman space, which is the closed subspace consisting of analytic functions. The Floquet transform is defined for $f\in A^{2}(\Omega)$ by
@@
The basic assumptions on the domain $\Omega$ and its periodic cells are as follows. However, we will pose one more assumption, related to the theory of elliptic functions, see $({\mathcal{A}})$ in the next section. Here and later, cl(A)(A) denotes the closure of the set AA.
Definition $1.$
@@
We emphasize that contrary to [10], we do not need to add assumptions on the smoothness of the boundary $\partial\Omega$. Anyway, the periodic domain $\Omega$ is always infinitely connected, due to assumption $(iii)$ of Definition $1.$
The following notation will be used throughout the paper. We write $C$, $C^{\prime}$, $\ldots$, (respectively, $C_{\eta}$, $C^{\prime}_{\eta}$, $\ldots$ etc.) for positive constants independent of the functions or variables in the given inequalities (resp. depending only on a parameter $\eta$ etc.), the values of which may vary from place to place. For $\in{\mathbb{R}}, [x][x$] denotes the largest integer not larger than $x$. We write ${\mathbb{Z}}=\{0,\pm 1,\pm 2,\ldots\}$. Subsets of the complex plane will often be described using the real planar coordinates, for example, rectangles in $\mathbb{C}$} are described as the sets $a,b]\times[c,d]$. If $x$ is a point in $\mathbb{C}$} or ${\mathbb{R}}^{n}$ and $r>0$, then $B(x,r)$ denotes the Euclidean ball with center $x$ and radius $r>0$. Moreover, cl(A)(A) denotes the closure of a set AA.
Given a domain $D\subset{\mathbb{C}}$, we denote by $\|f\|_{D}$ and $(\cdot|\cdot)_{D}$ the norm and inner product of $L^{2}(D)$, which is the $L^{2}$ space with respect to the (real) area measure. In general, the norm of a Banach space $X$ is denoted by $\|\cdot\|_{X}$. Given an interval $I\subset{\mathbb{R}}$ and a Banach space $X$ we denote by $L^{2}(I;X)$ the space of vector valued, Bochner-$L^{2}$-integrable functions on II with values in $X$, endowed with the norm
@@
Proof. We first observe that by Proposition 3.2.(ii)(ii), $\sf F$} maps the dense subspace $X$ of $A^{2}(\Omega)$ into ${\mathcal{H}}_{\eta}$. In view of the unitarity of $\sf F$}, this implies that ${\sf F}(A^{2}(\Omega))$ is contained in the closure of ${\mathcal{H}}_{\eta}$ in $L^{2}({\bf Q};L^{2}(\varpi))$, which equals $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$, by Lemma $3.$
@@ 5. Spectral band formula for Toeplitz operators with doubly periodic symbols
Moreover, the spectra of $T_{a,\eta}$ in $A_{\eta}^{2}(\varpi)$ and $T_{a,\eta_{k},\eta}$ in $A_{\eta_{k}}^{2}(\varpi)$ coincide (cf. Lemma $3.$ of [11]), which implies that, for a large enough $k$, $\lambda_{k}$ belongs to the resolvent set of $T_{a,\eta_{k},\eta}$, see ()(*). This eventually leads to a contradiction with $\lambda_{k}\in\sigma(T_{a,\eta_{k}})$ by using some quite standard resolvent estimates and the closeness of the two operators given by formula (5.8).
@@
To complete the proof of Theorem 5.1, it is thus sufficient to show that $\Sigma\subset\sigma_{\rm ess}(T_{a})$, i.e. every $\lambda$, which belongs to $\sigma(T_{a,\mu})$ for some $\mu\in{\bf Q}$, is a point in the essential spectrum of $T_{a}$. To this end, we will use the Weyl criterion as presented e.g. in Lemma $1.$ of [11], and thus we fix $\mu\in{\bf Q}$, $\lambda\in\sigma(T_{a,\mu})$, an arbitrary $\varepsilon>0$ and a near eigenfunction $g\in A_{\mu}^{2}(\varpi)$ with $\|g\|_{\varpi}=1$ such that

### v2

@@ Abstract.
Addresses:
1 Department of Mathematics and Statistics, University of Helsinki, P.O. Box 68, 00014 Helsinki, Finland
email jari.taskinen@helsinki.fi
2 Department of Mathematics and Statistics, University of Helsinki, P.O. Box 68, 00014 Helsinki, Finland
email zhan.zhang@helsinki.fi
@@
Given a domain $\Omega$ in the complex plane $\mathbb{C}$}, we denote by $L^{2}(\Omega)$ the usual Lebesgue-Hilbert space with respect to the (real) area measure $dA$ and by $A^{2}(\Omega)$ the corresponding Bergman space, which is the closed subspace consisting of analytic functions. The Floquet transform is defined for $f\in A^{2}(\Omega)$ by
@@
The basic assumptions on the domain $\Omega$ and its periodic cells are as follows. However, we will pose one more assumption, related to the theory of elliptic functions, see $({\mathcal{A}})$ in the next section. Here and later, cl(A)(A) denotes the closure of the set $A$.
Definition 1.1.
@@
We emphasize that contrary to [10], we do not need to add assumptions on the smoothness of the boundary $\partial\Omega$. Anyway, the periodic domain $\Omega$ is always infinitely connected, due to assumption $(iii)$ of Definition 1.1.
The following notation will be used throughout the paper. We write $C$, $C^{\prime}$, $\ldots$, (respectively, $C_{\eta}$, $C^{\prime}_{\eta}$, $\ldots$ etc.) for positive constants independent of the functions or variables in the given inequalities (resp. depending only on a parameter $\eta$ etc.), the values of which may vary from place to place. For $\in{\mathbb{R}}, [x][x$] denotes the largest integer not larger than $x$. We write ${\mathbb{Z}}=\{0,\pm 1,\pm 2,\ldots\}$. Subsets of the complex plane will often be described using the real planar coordinates, for example, rectangles in $\mathbb{C}$} are described as the sets $a,b]\times[c,d]$. If $x$ is a point in $\mathbb{C}$} or ${\mathbb{R}}^{n}$ and $r>0$, then $B(x,r)$ denotes the Euclidean ball with center $x$ and radius $r>0$. Moreover, cl(A)(A) denotes the closure of a set $A$.
Given a domain $D\subset{\mathbb{C}}$, we denote by $\|f\|_{D}$ and $(\cdot|\cdot)_{D}$ the norm and inner product of $L^{2}(D)$, which is the $L^{2}$ space with respect to the (real) area measure. In general, the norm of a Banach space $X$ is denoted by $\|\cdot\|_{X}$. Given an interval $I\subset{\mathbb{R}}$ and a Banach space $X$ we denote by $L^{2}(I;X)$ the space of vector valued, Bochner-$L^{2}$-integrable functions on $I$ with values in $X$, endowed with the norm
@@
Proof. We first observe that by Proposition 3.2.(ii)(ii), $\sf F$} maps the dense subspace $X$ of $A^{2}(\Omega)$ into ${\mathcal{H}}_{\eta}$. In view of the unitarity of $\sf F$}, this implies that ${\sf F}(A^{2}(\Omega))$ is contained in the closure of ${\mathcal{H}}_{\eta}$ in $L^{2}({\bf Q};L^{2}(\varpi))$, which equals $L^{2}({\bf Q};A_{\eta}^{2}(\varpi))$, by Lemma 3.3.
@@ 5. Spectral band formula for Toeplitz operators with doubly periodic symbols
Moreover, the spectra of $T_{a,\eta}$ in $A_{\eta}^{2}(\varpi)$ and $T_{a,\eta_{k},\eta}$ in $A_{\eta_{k}}^{2}(\varpi)$ coincide (cf. Lemma 3.3. of [11]), which implies that, for a large enough $k$, $\lambda_{k}$ belongs to the resolvent set of $T_{a,\eta_{k},\eta}$, see ()(*). This eventually leads to a contradiction with $\lambda_{k}\in\sigma(T_{a,\eta_{k}})$ by using some quite standard resolvent estimates and the closeness of the two operators given by formula (5.8).
@@
To complete the proof of Theorem 5.1, it is thus sufficient to show that $\Sigma\subset\sigma_{\rm ess}(T_{a})$, i.e. every $\lambda$, which belongs to $\sigma(T_{a,\mu})$ for some $\mu\in{\bf Q}$, is a point in the essential spectrum of $T_{a}$. To this end, we will use the Weyl criterion as presented e.g. in Lemma 1.1. of [11], and thus we fix $\mu\in{\bf Q}$, $\lambda\in\sigma(T_{a,\mu})$, an arbitrary $\varepsilon>0$ and a near eigenfunction $g\in A_{\mu}^{2}(\varpi)$ with $\|g\|_{\varpi}=1$ such that